OpenStudy (cookie_2046):

Good descriptive writing uses precise language such as? A. General adjectives B. Passive verbs C. Generic nouns D. Strong action verbs

OpenStudy (anonymous):

what do you think it should be @Cookie_2046

OpenStudy (cookie_2046):

a

OpenStudy (anonymous):

I think it might be a or d they both make sense

OpenStudy (cookie_2046):

So confusing :p

OpenStudy (cookie_2046):

@sourwing @vlz6000 @jcoury @animalfan1

OpenStudy (anonymous):

lol idk on this one but i think its A

OpenStudy (anonymous):

i need help[ if you help me ill give you a medal

OpenStudy (anonymous):

What do you need help with

OpenStudy (anonymous):

(original amount of juice + added juice) - juice that was drank = juice left in the jar (0.25 + 0.30 + 0.75) - 0.20 = 1.30 - 0.20 = 1.10 left in jar i need to know the property to this